OpenStudy (anonymous):

Hello, everyone. I'm not very sure about this question, so can you please give me a hint for this question? Thanks! Glucose is fed intravenous injection at a constant rate, k, into a patient's bloodstream. Once there, the glucose is removed at at rate proportional to the amount of glucose present. If R is the net rate at which the quantity, G, of glucose in the blood is increasing: a) Write a formula giving R as a function of G b) Sketch a graph of R against G. P.S. I cannot understand what R is.

OpenStudy (anonymous):

The amount of glucose=x\[\frac{dx}{dt}=k-cx\]

OpenStudy (anonymous):

Do you know anything about differential equations?

OpenStudy (anonymous):

Hello. Well, I do know about them, but this was supposed to be PreCalculus stuff.

OpenStudy (anonymous):

I think this requires at least knowledge of exponential growth (thus differentiation).

OpenStudy (wwe123):

Rate in = k. Rate out = mG (where m is the constant of proportionality; standard convention is to use k for that constant but we've already used that letter). Hence, R = k - mG.

OpenStudy (anonymous):

The only way to do it without calculus (*I think*) is to assume that\[x=Ae^{qt}\]

OpenStudy (wwe123):

a

OpenStudy (anonymous):

Oh no, re-read the question, I'm overcomplicating most probably.

OpenStudy (wwe123):

b

OpenStudy (anonymous):

Wait, I don't have to use anything like exponential growth?

OpenStudy (anonymous):

I was overcomplicating, it's only implicit in your question. It makes it much easier to solve, though.

OpenStudy (anonymous):

And it's not quite normal exponential growth: I guess the constant q is imaginary, so by Euler's formula is a trig function.